At the beginning of a walk, Roberto and Juanita are 4.4 miles apart. If they leave at the same time and walk in the same direction, Roberto overtakes Juanita in 11 hours. If they walk towards each other, they meet in hour. What are their speeds?
Roberto's speed is __ mph.
Juanita's speed is __ mph.

Answer by lwsshak3(11628)

At the beginning of a walk, Roberto and Juanita are 4.4 miles apart. If they leave at the same time and walk in the same direction, Roberto overtakes Juanita in 11 hours. If they walk towards each other, they meet in hour. What are their speeds?
Roberto's speed is __ mph.
Juanita's speed is __ mph.
***
Both walking in the same direction:
let x=distance Juanita travels
4.4+x=distance Roberto travels
speed=distance/travel time
Juanita's speed=x/11
Roberto's speed=(4.4+x)/11
...
Both walking toward each other:
let y=distance Juanita travels
4.4-y=distance Roberto travels
Juanita's speed=y/1=y
Roberto's speed=(4.4-y)/1=4.4-y
...
juanita's speed=x/11=y
x=11y
Roberto's speed=(4.4+x)/11)=4.4-y
=(4.4+11y)/11=4.4-y
=4.4+11y=48.4-11y
22y=44
y=2
4.4-y=2.2
Roberto's speed is 2.2 mph.
Juanita's speed is 2.0 mph.